OpenStudy (anonymous):

What are the real zeroes of x3 + 5x2 – 9x – 45?

Directrix (directrix):

You are looking for the real values of x that cause the given equation to equal zero. x^3 + 5x^2 – 9x – 45 = 0 Factor by grouping x^3 + 5x^2 – 9x – 45 = 0 x^2 ( x + 5) - 9 (x + 5) = 0 (x + 5) ( x^2 - 9) = 0

Directrix (directrix):

(x + 5) ( x^2 - 9) = 0 From here, your task is to contiinue the factorization by factoring (x^2 - 9). Then, use the Zero Product Property to solve for x. @jonathanpinon123